What is the average reading speed of Dante if he can read 500 pages in 3/4 days?
Answer by Theo(13342)

it depends on what the measurement is.

is it pages per day, or pages per hour, or pages per minute?

assuming it is pages per day, then do the following:

rate * time = quantity

rate is the average speed per day.
quantity is the number pages in 3/4 of a day.
time is 3/4 of a day.

since you don't know the rate per day, then fill in the other blanks to get:

rate * 3/4 = 500

solve for rate to get rate = 500 / (3/4) = 666.6666667 pages per day.

when rate is 666.6666667 pages per day and the time is 3/4 of a day, then the number of pages read is 666.6666667 * 3/4 = 500

this confirms the rate per day is correct.

you could also solve for hours by making the time equal to 3/4 * 24 = 18 hours.

you could also solve for minutes by making the time equal to 3/4 * 24 * 60 = 1080.

for example, if you're looking for the rate per minute, then the formula becomes:

rate * 1080 = 500

solve for rate to get:

rate = 500 / 1080 = .4629622963 pages per minute.

rate * time = quantity becomes:

.4629622963 * 1080 = quantity.

solve for quantity to get quantity = 500.

at a rate of .4629622963 pages per minute for 1080 minutes, 500 pages are read.

1080 minutes / 60 = 18 hours / 24 = .75 days = 3/4 of a day.